Find the radius of the circular section of the
sphere Ir—cl = 7 by the plane
r. (3i — j + 2k) = 2 ,r7 , where c = (— 1, 0, 1).
1
Expert's answer
2019-12-17T09:45:17-0500

Let 𝑆 be the sphere in 𝑅3 with center 𝑂(𝑥0, 𝑦0, 𝑧0) and radius 𝑟, and let 𝑃 be the plane with equation 𝐴𝑥 + 𝐵𝑦 + 𝐶𝑧 = 𝐷,


so that 𝑛⃗ = (𝐴, 𝐵, 𝐶) is a normal vector of 𝑃.

If 𝑃0 is an arbitrary point on 𝑃, the signed distance from the center of the sphere 𝑂 to the plane 𝑃 is

𝜌 = (𝑂 − 𝑃0)𝑛⃗/ |𝑛⃗|

= (𝐴𝑥0 + 𝐵𝑦0 + 𝐶𝑧0 − 𝐷)/(A2+B2+C2)1/2


The intersection 𝑆 ∩ 𝑃 is a circle if and only if −𝑟 < 𝜌 < 𝑟, and in that case, the circle has radius

𝑟c = "\sqrt{r^2-\rho^2}"  and

center 𝑐 = 𝑂 + 𝜌 ∙ 𝑛⃗/ |𝑛⃗ |

=(𝑥0, 𝑦0, 𝑧0) + 𝜌 ∙ (𝐴, 𝐵, 𝐶)/(A2+B2+C2 )1/2


so in our case,

|𝑟| =7 and 𝑂(𝑥0, 𝑦0, 𝑧0)=(-1,0,1) and P={3x - y + 2z = 2}


𝜌= (𝐴𝑥0 + 𝐵y0 + 𝐶𝑧0 −D)/(A2+B2+C2)1/2

= (3.(-1)+0+1.2-2)/(14)1/2

= -3/(14)1/2

=-0.80


𝑟c = (𝑟2 − 𝜌2 )1/2 = (49-0.64)1/2 =6.95,

hence the radius of this circular plane is 6.95.